# GTM: Simulating the World of Tools for AI Agents

### v1

@@ 4 Method
GTM is designed to simulate diverse tools at an affordable computational cost. We selected Qwen2.5-3B as our base model, striking a balance between capability and cost. GTM training consists of three key steps: (1) Tool generation, where we create a comprehensive collection of tool specifications; (2) Context Awareness Response Generation (CARG) pipeline, which generates contextually coherent responses by understanding tool-specific context and maintaining continuity across multiple tool invocations in multi-turn interactions. (3) Training, where we fine-tune the model to achieve our design goals.
@@ 5 Experiments and Analysis
For the first group of experiments, we utilized GTM models fine-tuned from Qwen2.5-1.5B and Qwen2.5-0.5B, with the API dataset split into training and validation sets. These GTM variants were compared against several common open-source models, including the Qwen2.5 series, InternLM2.5 series, and Llama3.1 series. To ensure objective evaluation, a Qwen2.5-72B model was employed as the judger to assess output quality across the three scenarios.
@@ 5.1 Evaluation of GTM’s Output Quality
The experimental results in Table 1 demonstrate that GTM models achieve significant advantages in generating high-quality tool responses. Specifically, GTM-0.5B (Avg: 83.9%) outperforms all baseline models of comparable scale, including Qwen2.5-0.5B-Instruct (45.1%), Llama-3.2-1B-Instruct (39.8%), and even larger models like InternLM2.5-7B (39.2%) and Llama-3.2-3B-Instruct (68.3%). GTM-1.5B (Avg: 89.4%) further surpasses stronger baselines such Qwen2.5-14B-Instruct (85.8%), despite having fewer parameters. A notable pattern across all models is their strong performance in single-turn scenarios (high ”All” scores), where correct API calling conventions can be learned without context management challenges. However, all model’s performance declines in multi-turn scenarios, where preserving contextual continuity becomes crucial—here, GTM models demonstrate superior consistency, with GTM-1.5B achieving a multi-turn ”All” score of 86.7%, significantly outperforming baselines of similar or larger scale. In error detection scenarios, GTM-0.5B and GTM-1.5B achieve ”All” scores of 94.5% and 86.1%, respectively, demonstrating exceptional helpfulness in identifying errors and generating appropriate error messages, which helps agents learn proper parameter formatting requirements in RL processes.
@@ 5.6 Ablation Study
To verify whether CARG’s generation process can inherently teach models contextual awareness, we trained a variant of GTM-0.5B using only single-turn and error generation data, excluding all multi-turn dialogues from the training set. As shown in Table 6, while this single-turn-only variant improved significantly over the base model (83.0% vs. 53.0% on multi-turn “All” score), it still falls short of the full GTM’s 86.7%. These results suggest that CARG can enable models to acquire contextual awareness even from single-turn data alone. However, to achieve optimal multi-turn performance, incorporating dedicated multi-turn training data remains essential.

### v2

@@ 4 Method
GTM is designed to simulate diverse tools at an affordable computational cost. We selected Qwen2.5-1.5B as our base model, striking a balance between capability and cost. GTM training consists of three key steps: (1) Tool generation, where we create a comprehensive collection of tool specifications; (2) Context Awareness Response Generation (CARG) pipeline, which generates contextually coherent responses by understanding tool-specific context and maintaining continuity across multiple tool invocations in multi-turn interactions. (3) Training, where we fine-tune the model to achieve our design goals.
@@ 5 Experiments and Analysis
For the first group of experiments, we utilized GTM models fine-tuned from Qwen2.5-1.5B, with the API dataset split into training and validation sets. These GTM variants were compared against several common open-source models, including the Qwen2.5 series, InternLM2.5 series, and Llama3.1 series. To ensure objective evaluation, a Qwen2.5-72B model was employed as the judger to assess output quality across the three scenarios.
@@ 5.1 Evaluation of GTM’s Output Quality
The experimental results in Table 1 demonstrate that GTM models achieve significant advantages in generating high-quality tool responses. Specifically, GTM-1.5B (Avg: 89.4%) further surpasses stronger baselines such Qwen2.5-14B-Instruct (85.8%), despite having fewer parameters. A notable pattern across all models is their strong performance in single-turn scenarios (high ”All” scores), where correct API calling conventions can be learned without context management challenges. However, all model’s performance declines in multi-turn scenarios, where preserving contextual continuity becomes crucial, here, GTM models demonstrate superior consistency, with GTM-1.5B achieving a multi-turn ”All” score of 86.7%, significantly outperforming baselines of similar or larger scale.
@@ 5.6 Ablation Study
To verify whether CARG’s generation process can inherently teach models contextual awareness, we trained a variant of GTM-1.5B using only single-turn and error generation data, excluding all multi-turn dialogues from the training set. As shown in Table 6, while this single-turn-only variant improved significantly over the base model (83.0% vs. 53.0% on multi-turn “All” score), it still falls short of the full GTM’s 86.7%. These results suggest that CARG can enable models to acquire contextual awareness even from single-turn data alone. However, to achieve optimal multi-turn performance, incorporating dedicated multi-turn training data remains essential.